### v1

# An Accurate Efficient Analytic Model of Fidelity under Depolarizing Noise oriented to Large Scale Quantum System Design

Abstract
Fidelity is one of the most valuable and commonly used metrics for assessing the performance of quantum circuits on error-prone quantum processors. Several approaches have been proposed to estimate circuit fidelity without the need of executing it on quantum hardware, but they often face limitations in scalability or accuracy. In this work, we present a comprehensive theoretical framework to predict the fidelity of quantum circuits under depolarizing noise. Building on theoretical results, we propose an efficient fidelity estimation algorithm based on device calibration data. The method is thoroughly validated through simulation and execution on real hardware, demonstrating improved accuracy compared to state-of-the-art alternatives. The proposed approach provides a scalable and practical tool for benchmarking quantum hardware, comparing quantum software techniques such as compilation methods, obtaining computation bounds for quantum systems, and guiding hardware design decisions, making it a critical resource for the development and evaluation of quantum computing technologies.
keywords:
Quantum Information, Quantum Computing, Depolarizing Noise, Quantum Fidelity
1 Introduction
Quantum errors are one of the most significant challenges in current quantum computers [1, 2], limiting the reliable execution of quantum algorithms. Quantum systems suffer from various types of errors, including thermal errors [3, 4], crosstalk errors [5, 6], readout errors [7], and operational errors, which occur when executing quantum gates. Quantum error correction [8, 9, 10, 11, 12, 13] and quantum error mitigation [14, 15, 16] techniques have been proposed to suppress or mitigate those errors, increasing the reliability of quantum computations and enabling practical use of quantum devices for tasks such as optimization [17], cryptography [18], and simulation [19].
Quantum fidelity [20, 21] is a measure of the similarity between two quantum states and hence serves as an indicator of either the reliability of a computation or the presence of excessive errors. It is calculated using the density matrices representing the quantum states ($\rho$ and $\sigma$), as defined in Equation (1).
$F(\rho,\sigma):=||\sqrt{\rho}\sqrt{\sigma}||_{1}^{2}=\left(tr\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}\right)^{2}$ (1)
If at least one of the two states ($\rho$ or $\sigma$) is pure, the fidelity can be expressed as the trace of the product of the pure state’s density matrix and the other state, whether pure or mixed (i.e., $F(\rho,\sigma)=tr(\sigma\rho)$). Fidelity exhibits properties such as unitary invariance and multiplicativity under tensor product (Equations (2) and (3)) [21], both of which are essential for the analysis presented in this work.
$F(\rho,\sigma)=F(U\rho U^{\dagger},U\sigma U^{\dagger})$ (2)
$F(\rho_{1}\otimes\rho_{2},\sigma_{1}\otimes\sigma_{2})=F(\rho_{1},\sigma_{1})\cdot F(\rho_{2},\sigma_{2})$ (3)
Estimating the fidelity of quantum circuits without executing them on quantum hardware is crucial to advance in the design of quantum processors. A reliable fidelity estimation method enables researchers to compare quantum systems, optimize compilation strategies, and predict performance across different hardware platforms. As the number of qubits increases, the estimation method must remain computationally efficient to scale with the growing complexity of quantum circuits.
Various techniques have been proposed to estimate the quantum fidelity of a given quantum circuit when executed on a specific quantum processor. These approaches include simulation-based methods, analytical models, and machine-learning strategies.
Simulating quantum circuits [22, 23, 24, 25] enables adding noise to the quantum state, a process known as statistical fault injection [26]. The fidelity of the computation can be assessed by comparing a simulated noisy state with its noiseless counterpart. However, the simulation of quantum states is computationally expensive, with complexity growing exponentially with the number of qubits, restricting its application to small-scale circuits [27]. The accuracy of the estimation will depend on the noise model used in the simulation, as well as on the precision of the simulation itself, which may introduce errors arising from approximations (e.g., limiting the dimensionality of tensors in a tensor network simulation [28]).
Analytical approaches for fidelity estimation [29, 30] typically rely on gate error rates obtained from device calibration, usually through randomized benchmarking [31, 32], without assuming any error model in particular. An example of such methods is the Probability of Successful Trial (PST) [29], also known as Estimated Success Probability (ESP), a commonly used metric for estimating the fidelity of a quantum computation by multiplying the individual gate fidelities ($F_{g_{i}}$) and measurement fidelities ($F_{m_{i}}$) of the circuit (Equation (4)). However, ESP does not account for how errors propagate through the system or specify which qubits are affected by the errors, obtaining a circuit fidelity that does not depend on the circuit structure but just on the number of gates and their respective fidelities.
$ESP=\prod_{i=1}^{N_{gates}}F_{g_{i}}\cdot\prod_{i=1}^{N_{meas}}F_{m_{i}}$ (4)
The Quantum Vulnerability Analysis (QVA) method [30], like ESP, uses individual gate fidelities as its basis. However, it introduces a hyperparameter ($w\in[0,1]$) that quantifies the degree of cross-error introduced by a two-qubit gate, requiring fine-tuning for each specific system. Additionally, unlike ESP, QVA accounts for the fidelity of each two-qubit gate twice (once for each qubit involved), resulting in consistently lower fidelity estimates compared to ESP. Machine-learning-based techniques [33, 34, 35] have also been proposed to estimate circuit fidelity. These approaches often target specific platforms, requiring extensive training data and retraining for different devices.
To better understand the strengths and limitations of existing fidelity estimation methods, we compare their predictions against experimental results from a real quantum processor.
Figure 1 compares the fidelity predictions of statistical fault injection (using Qiskit’s simulator [36]), ESP, and QVA for quantum circuits ranging from 2 to 8 qubits, against the success rate of the same circuits executed on IBM Q Kyiv (Eagle r3 processor [37]). The process from obtaining the circuits to computing the success rate after readout mitigation is explained in Section 4.2. Figure 1 shows how both Qiskit’s simulation and ESP inaccurately predict the fidelity observed in real quantum processors, consistently overestimating it. Conversely, QVA’s performance highly depends on its hyperparameter ($w$), with fidelity predictions varying significantly, up to 0.8, between the extreme values of $w=0$ and $w=1$ (QVA max and QVA min in Figure 1). This introduces the additional challenge of determining the optimal $w$, which the original work [30] addressed by training a machine learning model for this purpose. In our analysis, we use $w=0.5$ (depicted as red points in Figure 1) as it provides a balanced prediction (see box plot in Figure 1), while also reporting results for $w=0$ and $w=1$. In the original work [30] the authors identify $w=0$ as the optimal value for large circuits.
Figure 1: Predicted fidelity using various state-of-the-art prediction methods alongside the measured success rate from real hardware. For the QVA method, vertical lines indicate the range of solutions corresponding to different values of the $w$ hyperparameter. The size of the points is proportional to the number of qubits in the circuit (between 2 and 8). Boxplots display the distance between the predicted fidelity ($fid_{pred}$) and the measured success rate ($sr_{true}$), quantified as $fid_{pred}-sr_{true}$, providing insight into the accuracy and variability of each model’s predictions.
These observations highlight the need for a more accurate fidelity estimation model that, like QVA and ESP, is scalable in terms of computational cost, allowing to increase the number of qubits and gates in the circuit. However, it is essential to minimize the dependence on hyperparameters that can excessively influence predictions, as this variability can compromise the reliability and consistency of the model’s outcomes.
In this work, we develop a theoretical model for estimating the fidelity under depolarizing noise (Equation (5)), in which each quantum gate introduces depolarization to the involved qubits, a common assumption regarding operational errors [38, 39, 40, 41, 42, 2]. By examining how depolarizing noise impacts circuit fidelity, we lay the groundwork for a scalable fidelity estimation algorithm rooted in theoretical principles.
$\rho\rightarrow\mathcal{E}(\rho)=(1-p)\rho+\frac{p}{3}(X\rho X+Y\rho Y+Z\rho Z)=(1-p)\rho+\frac{p}{d}I_{d}$ (5)
We will primarily use the last expression in Equation (5) to represent the depolarizing channel, where $d$ is the dimensionality of the quantum system being depolarized, $I_{d}/d$ is the maximally mixed state with the same dimensions ($d$) as $\rho$ (the quantum state described by its density matrix), and $p$ is the depolarization factor (or depolarization probability). This formulation provides a compact and general framework to model depolarization effects across systems of varying dimensions.
2 Results
2.1 Theoretical Modeling of the Fidelity Loss
In this section, we develop and prove the theorems that establish the foundation for a theoretical analysis of how the depolarizing noise channel influences the fidelity of quantum states.
To understand how depolarizing noise affects quantum circuits, we begin by exploring the evolution of a quantum state under the influence of repeated depolarizing channels. Theorem 1 provides a formal expression for how a quantum state, described by its density matrix $\rho$, evolves after $n$ depolarizing channels have been applied, each introducing a depolarizing factor $p$.
Theorem 1 (Consecutive depolarizing channels).
After applying $n$ consecutive depolarizing channels with depolarizing factor $p$ to an initial pure quantum state $\rho$, the resulting quantum state $\mathcal{E}^{n}(\rho)$ can be expressed as:
$\mathcal{E}^{n}(\rho)=(1-p)^{n}\rho+\frac{1-(1-p)^{n}}{d}I_{d}$ (6)
Proof.
By induction.
Let $P(n)$ be the statement that $\mathcal{E}^{n}(\rho)=(1-p)^{n}\rho+\frac{1-(1-p)^{n}}{d}I_{d}$.
Base case: For $n=1$ we have:
$\mathcal{E}^{1}(\rho)=(1-p)^{1}\rho+\frac{1-(1-p)^{1}}{d}I_{d}=(1-p)\rho+\frac{p}{d}I_{d}$ (7)
This is true by the definition of depolarizing channel given in Equation(5).
Inductive hypothesis: Assume $P(k)$ is correct for some positive integer $k$, which means $\mathcal{E}^{k}(\rho)=(1-p)^{k}\rho+\frac{1-(1-p)^{k}}{d}I_{d}$.
Inductive step: We now show that $P(k+1)$ is correct.
$\mathcal{E}^{k+1}(\rho)$ $=\mathcal{E}\left(\mathcal{E}^{k}(\rho)\right)$ (8)
$=(1-p)\left((1-p)^{k}\rho+\frac{1-(1-p)^{k}}{d}I_{d}\right)+\frac{p}{d}I_{d}$ (9)
$=(1-p)^{k+1}\rho+\frac{(1-p)-(1-p)^{k+1}}{d}I_{d}+\frac{p}{d}I_{d}$ (10)
$=(1-p)^{k+1}\rho+\frac{1-(1-p)^{k+1}}{d}I_{d}$ (11)
Following this, Theorem 2 quantifies the impact of these repeated depolarizations on the fidelity of the quantum state.
Theorem 2 (Fidelity after $n$ depolarizations).
The fidelity between a pure quantum state, $\rho$, and the same quantum state after $n$ depolarizing channels, $\mathcal{E}^{n}(\rho)$, is given by:
$F(\rho,\mathcal{E}^{n}(\rho))=(1-p)^{n}+\frac{1-(1-p)^{n}}{d}$ (12)
Proof.
$F\left(\rho,\mathcal{E}^{n}(\rho)\right)$ $=tr\left(\rho\cdot\mathcal{E}^{n}(\rho)\right)$ (13)
$=tr\left(\rho\cdot\left((1-p)^{n}\rho+\frac{1-(1-p)^{n}}{d}I_{d}\right)\right)$ (14)
$=tr\left((1-p)^{n}\rho^{2}\right)+tr\left(\frac{1-(1-p)^{n}}{d}\rho\right)$ (15)
$=(1-p)^{n}tr\left(\rho^{2}\right)+\frac{1-(1-p)^{n}}{d}tr\left(\rho\right)$ (16)
$=(1-p)^{n}+\frac{1-(1-p)^{n}}{d}$ (17)
To further explore the impact of depolarization, Corollary 1 analyses the fidelity loss after a single depolarization, considering the effects of repeated depolarizations over time. This result will be used to quantify each gate error’s contribution when estimating the fidelity of quantum circuits.
Corollary 1 (Fidelity loss).
The change in fidelity from a single depolarizing channel $\mathcal{E}^{n+1}(\rho)=\mathcal{E}(\mathcal{E}^{n}(\rho))$ applied to an already depolarized state $\mathcal{E}^{n}(\rho)$ (previously pure state $\rho$) is:
$F\left(\rho,\mathcal{E}^{n+1}(\rho)\right)=(1-p)F\left(\rho,\mathcal{E}^{n}(\rho)\right)+\frac{p}{d}$ (18)
Proof.
$F\left(\rho,\mathcal{E}^{n+1}(\rho)\right)$ $=tr\left(\rho\cdot\mathcal{E}^{n+1}(\rho)\right)$ (19)
$=tr\left(\rho\cdot\left((1-p)\mathcal{E}^{n}(\rho)+\frac{p}{d}I_{d}\right)(\rho)\right)$ (20)
$=(1-p)tr\left(\rho\cdot\mathcal{E}^{n}(\rho)\right)+\frac{p}{d}tr(\rho I_{d})$ (21)
$=(1-p)F\left(\rho,\mathcal{E}^{n}(\rho)\right)+\frac{p}{d}$ (22)
Theorem 3 extends the previous results by focusing on a composite quantum state. We assume that the depolarizing error acting jointly on both subsystems can be decomposed into independent errors on each subsystem, leveraging the multiplicativity under tensor product property of fidelity (Equation (3)). This enables us to track the fidelity of individual qubits in a multi-qubit system, which will be the foundation of the proposed fidelity estimation algorithm, described in Section 4.1.
Theorem 3 (Fidelity loss in a composite state).
Let $\rho=\rho_{A}\otimes\rho_{B}$ be a pure, product state, and $\rho^{\prime}=\mathcal{E}^{i}(\rho_{A})\otimes\mathcal{E}^{j}(\rho_{B})$ be the same product state after each component has undergone its respective depolarization process. The fidelity loss that a new depolarizing channel $\mathcal{E}(\rho^{\prime})=\mathcal{E}\left(\mathcal{E}^{i}(\rho_{A})\otimes\mathcal{E}^{j}(\rho_{B})\right)$ applies to each previously depolarized quantum state is:
$F\left(\rho_{A},\mathcal{E}(\rho^{\prime})_{A}\right)=\sqrt{1-p}\cdot F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)+\eta$ (23)
$F\left(\rho_{B},\mathcal{E}(\rho^{\prime})_{B}\right)=\sqrt{1-p}\cdot F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)+\eta$ (24)
where $\mathcal{E}(\rho^{\prime})_{A}$ and $\mathcal{E}(\rho^{\prime})_{B}$ are the subsystems $A$ and $B$ in the composite state $\mathcal{E}(\rho^{\prime})$, with dimensionality $d_{AB}$, and:
$\eta=$ $\frac{1}{2}\bigg{(}-\sqrt{1-p}\left(F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)+F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)\right)$
$+\sqrt{(1-p)\left(F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)+F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)\right)^{2}+\frac{4p}{d_{AB}}}\quad\bigg{)}$ (25)
Proof.
$F(\rho,\mathcal{E}(\rho^{\prime}))$ $=(1-p)\cdot F\left(\rho,\rho^{\prime}\right)+\frac{p}{d_{AB}}$ (26)
$=(1-p)\cdot F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)\cdot F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)+\frac{p}{d_{AB}}$ (27)
$=\sqrt{1-p}\cdot F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)\cdot\sqrt{1-p}\cdot F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)+\frac{p}{d_{AB}}$ (28)
exists $\eta>0$ s.t.
$=\left(\sqrt{1-p}\cdot F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)+\eta\right)\cdot\left(\sqrt{1-p}\cdot F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)+\eta\right)$ (29)
$=F\left(\rho_{A},\mathcal{E}(\rho^{\prime})_{A}\right)\cdot F\left(\rho_{B},\mathcal{E}(\rho^{\prime})_{B}\right)$ (30)
From Equations (28) and (29) we have:
$\eta^{2}+\eta\sqrt{1-p}\left(F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)+F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)\right)-\frac{p}{d_{AB}}=0$ (31)
$\eta=$ $\frac{1}{2}\bigg{(}-\sqrt{1-p}\left(F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)+F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)\right)$
±$\pm\sqrt{(1-p)\left(F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)+F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)\right)^{2}+\frac{4p}{d_{AB}}}\quad\bigg{)}$ (32)
Since $\sqrt{1-p}\left(F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)+F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)\right)>0$, and $\eta>0$ we can discard the negative solution, resulting in
$\eta=$ $\frac{1}{2}\bigg{(}-\sqrt{1-p}\left(F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)+F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)\right)$
$+\sqrt{(1-p)\left(F\left(\rho_{A},\mathcal{E}^{i}(\rho_{A})\right)+F\left(\rho_{B},\mathcal{E}^{j}(\rho_{B})\right)\right)^{2}+\frac{4p}{d_{AB}}}\quad\bigg{)}$ (33)
Which guarantees $\eta\geq 0$. ∎
Finally, Theorem 4 addresses the effect of depolarization on the fidelity of entangled qubits. This theorem explores the fidelity loss in a qubit entangled with another quantum state when a depolarizing channel is applied only to a subset of the entangled system.
Theorem 4 (Fidelity loss in an entangled system).
Let $\rho_{AB}$ be a composite state in the $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ space. Suppose a depolarizing channel $\mathcal{E}$ is applied only to subsystem $A$. Let $\mathcal{E}_{A}(\rho_{AB})=(1-p)\rho_{AB}+\frac{p}{d_{A}}\left(I_{d_{A}}\otimes tr_{A}(\rho_{AB})\right)$ be the resulting state.
The fidelity of subsystem $A$ between $\rho_{AB}$ and $\mathcal{E}_{A}(\rho_{AB})$ is bounded by:
$(1-p)\leq F\left(\rho_{AB},\mathcal{E}_{A}(\rho_{AB})\right)\leq(1-p)+\frac{p}{d_{A}}$
Proof.
$F\left(\rho_{AB},\mathcal{E}_{A}(\rho_{AB})\right)$ $=tr\left(\rho_{AB}\cdot\mathcal{E}_{A}(\rho_{AB})\right)$ (34)
$=tr\left(\rho_{AB}\cdot\left[(1-p)\rho_{AB}+\frac{p}{d_{A}}\left(I_{d_{A}}\otimes tr_{A}(\rho_{AB})\right)\right]\right)$ (35)
$=(1-p)tr(\rho_{AB}^{2})+\frac{p}{d_{A}}tr\left(\rho_{AB}\cdot\left(I_{d_{A}}\otimes tr_{A}(\rho_{AB})\right)\right)$ (36)
$=(1-p)+\frac{p}{d_{A}}tr\left(tr_{B}(\rho_{AB})\right)tr\left(\left(tr_{A}(\rho_{AB})\right)^{2}\right)$ (37)
$=(1-p)+\frac{p}{d_{A}}tr\left(\left(tr_{A}(\rho_{AB})\right)^{2}\right)$ (38)
If $\rho_{AB}$ is a pure product (i.e., separable) state, then $\rho_{B}=tr_{A}(\rho_{AB})$ is also a pure state and $tr\left(\left(tr_{A}(\rho_{AB})\right)^{2}\right)=1$. If $\rho_{AB}$ is entangled (and thus not a product state), then $\rho_{B}$ is not pure, and $tr\left(\left(tr_{A}(\rho_{AB})\right)^{2}\right)<1$.
The exact value of $tr\left(\left(tr_{A}(\rho_{AB})\right)^{2}\right)$ lies between 0 and 1 and depends on the state $\rho_{AB}$, which can only be known through the simulation of the whole quantum state. From these values, we derive upper and lower bounds for the loss of fidelity:
$(1-p)\leq F(\rho_{AB},\mathcal{E}_{A}(\rho_{AB}))\leq(1-p)+\frac{p}{d_{A}}$
Theorem 4 shows that the fidelity is upper bounded by the results found in Corollary 1, obtaining a lower fidelity (higher fidelity loss) whenever the qubit being depolarized is entangled with other systems.
By presenting these theorems, a rigorous framework for understanding and quantifying the effects of depolarizing noise on quantum states and circuits is built, providing the foundation for the proposed fidelity estimation algorithm (see Section 4.1), which, unlike ESP, QVA, or other analytical estimation techniques, is based on the theoretical impact of quantum errors on the quantum states, and not only on the device calibration data.
The algorithm developed using this theoretical framework avoids simulating the quantum circuit, making it inherently scalable to larger systems. This scalability comes at the cost of introducing a hyperparameter ($p_{ent}\in[0,1]$), which accounts for the entanglement level of the system to remain consistent with the findings in Theorem 4. By adjusting $p_{ent}$, the model can effectively approximate the fidelity bounds for varying levels of entanglement without requiring computationally expensive simulations, obtaining lower and upper bounds on circuit fidelity for $p_{ent}=1$ and $p_{ent}=0$ respectively.
Figure 2 presents a comparison between the fidelities predicted by our theoretical model and those obtained from statistical fault injection simulations. The results span 93 quantum circuits, varying in size from 2 to 8 qubits and containing up to 5000 gates. Circuits are compiled to match with the restrictions of the superconducting processor used in latter experiments (see Section 4.2).
As expected, the fidelity predictions align closely with the simulated values, though slight discrepancies arise due to the theoretical upper and lower bounds employed in our model. To better illustrate these discrepancies, Figure 2 also shows the absolute difference between the predicted (with $p_{ent}=0.5$) and simulated fidelities. The observed deviations remain within a reasonable range (for all data points the absoulte fidelity difference remains under 0.07), confirming that our approach provides a robust approximation of circuit fidelity while being significantly more efficient than direct simulation.
Figure 2: Comparison of predicted versus simulated fidelities with single-qubit gates depolarization of $p_{1}=10^{-3}$ and two-qubit gates depolarization of $p_{2}=5\cdot 10^{-3}$. In the left plot, each point illustrates the predicted ($x$-axis) against the simulated fidelity ($y$-axis), with the diagonal line ($x=y$) indicating perfect predictions. Horizontal lines indicate the range of predicted fidelities arising from the variability of the hyperparameter $p_{ent}$. The right plot shows the difference on fidelity ($fid_{pred}-fid_{simul}$).
2.2 Quantum Circuit Fidelity Estimation
Having validated the model through simulation, we now assess its performance against real quantum hardware. Using the same set of compiled circuits, we compare the predicted fidelities to the actual success rates (i.e., the number of correct measurements divided by the total number of executions) measured on an IBM Q superconducting processor [37].
The fidelity was estimated using our model alongside other methods, including Qiskit simulation (using noise models associated with the used processor), ESP, and QVA. We compared these predictions to the success rates obtained from executions on quantum hardware after readout error mitigation. Additionally, a version of each model incorporating coherence errors ($T_{1}$ and $T_{2}$), detailed in Section 4.1, were also considered to provide a more comprehensive assessment of a real execution. Since quantum hardware is subject to additional sources of error beyond those captured by the considered models, it is not possible to achieve an accurate prediction by considering only gate errors, making the inclusion of $T_{1}$ and $T_{2}$ essential for a robust validation against quantum hardware.
The execution flow and the detailed implementation of the proposed fidelity estimation algorithm are presented in Section 4.
Figure 3 illustrates the measured success rate in the quantum device (IBM Q Eagle processor) after readout error mitigation (on the $y$-axis) versus the predicted fidelity (on the $x$-axis) for each estimation method. In this plot, a perfect prediction would place the data point along the diagonal. For the QVA and proposed models, horizontal lines indicate the influence of the hyperparameters in each algorithm (i.e., $w$ for QVA and $p_{ent}$ for the proposed model), with the reported point representing the average value between the extremes of these parameters. Most gate-error models tend to estimate a higher fidelity than the actual observation on the real device, which is expected since the models do not account for all types of errors present in actual hardware (e.g., coherence or crosstalk errors). This effect is evident when incorporating coherence errors into the estimation models, as it demonstrates a notable improvement in aligning predictions with the observed success rates for all models except for QVA, where it leads to a worse prediction, estimating a lower fidelity than the observed one.
Figure 3: Comparison of predicted versus measured fidelities for all estimation models. Each scatter plot illustrates the predicted fidelity ($x$-axis) against the measured success rate ($y$-axis), with the diagonal line ($x=y$) indicating perfect predictions. For the QVA and the proposed model, horizontal lines indicate the range of predicted fidelities arising from the variability of their hyperparameters $w$ and $p_{ent}$ respectively.
Interestingly, QVA sometimes predicts a lower fidelity than the observed success rate despite not including coherence or crosstalk errors in its model. This discrepancy arises because QVA decreases the global fidelity twice for each two-qubit gate, doubling the reduction in fidelity compared to the actual gate error. When coherence errors are incorporated into the QVA model, the predicted fidelities decrease even further, as shown in the lower row of Figure 3. Moreover, the long horizontal lines for the QVA in Figure 3 highlight the significant influence that the hyperparameter for the cross-error ($w$) has on its predictions.
Surprisingly, we observe a significant discrepancy between Qiskit’s fidelity estimation, obtained through full circuit simulation, and the measured fidelity on real hardware. Given that the simulation incorporates noise models inferred from the quantum device, we expected its predictions to be much closer to the actual fidelity. This deviation suggests potential inaccuracies in the extracted noise models.
Figure 4 presents the difference between each circuit’s predicted and observed fidelity.
Figure 4: Comparison of the difference ($fid_{pred}-sr_{true}$) between predicted fidelity ($fid_{pred}$) and measured success rate ($sr_{true}$), plotted against the number of gates in each circuit for all fidelity estimation approaches.
Across all estimation techniques, we observe that, when only taking into account gate errors (i.e., upper row), for circuits with either very few or very many gates, the predictions tend to be relatively accurate. The explanation lies in the simplicity of estimating fidelity in these cases. For circuits with a very low number of gates, the fidelity is typically close to one, and fewer interactions among qubits reduce cross-talk errors, which are not considered by none of the evaluated models. Additionally, coherence noise effects (which are not accounted for in ESP and QVA) are less significant due to the shorter execution times of such circuits. Conversely, for circuits with a high number of gates, the fidelity generally approaches zero, causing the absolute difference between the predicted and real success rates to decrease.
To assess the accuracy of the proposed prediction models, we compute several regression metrics typically used to evaluate the performance of predictive models. These include the Mean Absolute Error (MAE) [43], which measures the average absolute difference between predicted and actual values; the Mean Squared Error (MSE) [44], which measures the average of the squares of the errors, that is, the average squared difference between the estimated values and the actual value; and the $R^{2}$ Score [45], which quantifies the proportion of variance in the observed data that is predictable from the model. Finally, the Pearson Correlation [46] coefficient is used to assess the linear relationship between the predicted and actual values, where a higher value indicates a stronger linear relationship. Additionally, the results with the cross-error parameter set to zero ($w=0$) in the QVA algorithm are also included since, in the analysis shown in [30], for most large circuits (over 100 two-qubit gates), the optimal $w$ was close to 0.
Mean Absolute Mean Squared $R^{2}$ Score $\uparrow$ Pearson
Error (MAE) $\downarrow$ Error (MSE) $\downarrow$ Correlation $\uparrow$
Qiskit 0.422 0.273 -0.831 0.606
Qiskit +$T_{1,2}$ 0.094 0.024 0.843 0.927
ESP 0.229 0.098 0.341 0.827
ESP +$T_{1,2}$ 0.080 0.016 0.891 0.945
QVA 0.134 0.034 0.769 0.903
QVA +$T_{1,2}$ 0.161 0.051 0.658 0.898
QVAw=0 0.126 0.043 0.715 0.895
QVAw=0 +$T_{1,2}$ 0.088 0.015 0.899 0.951
This Work 0.210 0.084 0.434 0.849
This Work +$T_{1,2}$ 0.067 0.008 0.944 0.976
Table 1: Regression metrics (Mean Absolute Error, Mean Squared Error, $R^{2}$ Score, and Pearson Correlation) for each fidelity estimation method: Qiskit simulation, ESP, QVA, and the proposed model, all with and without $T_{1,2}$. The best value for each metric is highlighted in bold, showcasing the relative performance of the approaches in predicting circuit fidelity with respect to the measured success rate on real hardware. Arrows indicate the preferred direction for each metric: for metrics where a higher value is better, the arrow points upwards ($\uparrow$), while for those where a lower value is more desirable, the arrow points downwards ($\downarrow$).
The results presented in Table 1 highlight the performance of the proposed model (including $T_{1}$ and $T_{2}$), which consistently outperforms all other approaches across the selected metrics. It achieves the lowest Mean Absolute Error, improving between 15.63% and 84.07% over other techniques; the lowest Mean Squared Error, with improvements ranging from 44.14% to 96.92%; the highest $R^{2}$ score, with gains of 4.96% to 213.54%; and the highest Pearson correlation, improving by 2.59% to 61.05%. The results including $T_{1,2}$ are a significant improvement over all models, demonstrating the importance of incorporating coherence noise into the depolarizing model. This suggests that including additional noise factors can significantly enhance the model’s accuracy in predicting the fidelity of quantum circuits executed on a real device, making it a highly promising method for future quantum error modeling and prediction.
Since the QVA model already predicts a lower fidelity than the observed success rate even without accounting for $T_{1}$ and $T_{2}$ errors, incorporating coherence errors further worsens its predictions. This underscores the need for a more robust theoretical foundation.
To ensure the robustness of our results, we repeated the experiments multiple times, conducting a total of six independent runs. In all cases, the proposed model consistently achieved the best performance across different calibration instances and quantum processors. Supplementary Figure 8 provides a detailed comparison of the prediction differences observed across all estimation methods, as well as the obtained $R^{2}$ scores for each experiment.
3 Discussion
The performance differences between the considered fidelity estimation models reveal key insights into their underlying assumptions and applicability. The QVA model, despite showing competitive performance in regression metrics, predicts lower fidelities than those observed in real hardware for some circuits. This behaviour, combined with its exclusion of significant error sources such as coherence noise, suggests incomplete theoretical foundations. While the QVA model may provide reasonable predictions in some contexts, its tendency to deviate from hardware-observed outcomes raises questions about its reliability in diverse scenarios. It is, therefore, crucial not only to evaluate models based on final metrics but also to examine their assumptions and mechanisms, ensuring they align with the complexities of real devices.
Figure 5: Required two-qubit depolarizing error rates ($p_{2}$) for achieving a circuit fidelity above 0.99 as the number of qubits increases, evaluated for various quantum circuits; QFT, Quantum Volume, Cuccaro Adder, Draper Adder, and GHZ State, obtained from Qiskit’s circuit library [36]. Each curve corresponds to a fixed single-qubit depolarizing error rate ($p_{1}$), with $p_{1}=0$, $10^{-7}$, and $10^{-6}$. The simulation is terminated when $p_{2}$ falls below $p_{1}$, indicating that achieving fidelity $>0.99$ is not feasible for that number of qubits with the specified $p_{1}$. Coherence (i.e., $T_{1}$ and $T_{2}$) errors are excluded from this analysis, as the focus is solely on evaluating the impact of operational errors for design-oriented considerations.
Unlike simulations requiring extensive computational resources, the scalable depolarizing model proposed in this work can efficiently and accurately predict circuit fidelities, opening new avenues for exploration. For instance, it enables comparative studies of compilation strategies seeking to maximize computation fidelity, as showcased in existing literature [30]. Additionally, the model facilitates the exploration of computation bounds for varying architectural parameters. By adjusting single- and two-qubit depolarizing factors ($p_{1}$ and $p_{2}$), the model can determine the conditions necessary to achieve high-fidelity computation. Figure 5 serves as an example. It illustrates, for various quantum circuits, the required $p_{2}$ values for achieving a circuit fidelity above 0.99 as the number of qubits increases for fixed $p_{1}$ values.
Such evaluations provide valuable insights for guiding quantum processor design. By establishing theoretical bounds on the required single- and two-qubit gate error rates, hardware developers can set performance targets before fabrication. This enables a more informed approach to optimizing hardware parameters, ensuring that the resulting quantum processors meet the fidelity requirements for practical computation.
Similarly, Figure 6 showcases the fidelity trends of Shor’s algorithm [18] with 18 qubits for different combinations of $p_{1}$ and $p_{2}$. These results underscore the model’s potential to guide architectural design and optimization, making it a powerful tool for advancing quantum computation.
These types of analysis are only possible thanks to the proposed estimation method’s accuracy and scalability. If the model were inaccurate, it would lead to misleading predictions, potentially guiding hardware design and algorithmic decisions in the wrong direction. Additionally, without it being scalable, it would be unfeasible to thoroughly explore the full range of architectural parameters or extend the analysis to a larger number of qubits. By enabling precise and efficient fidelity estimations, the proposed approach allows for an in-depth evaluation of different quantum computing scenarios, supporting the development of more reliable and optimized quantum hardware and software.
Figure 6: Estimated circuit fidelity for Shor’s algorithm [18], containing 18 qubits, as a function of single-qubit ($p_{1}$) and two-qubit ($p_{2}$) gate depolarizing factors. The values of $p_{1}$ range from 0 to $10^{-5}$, while $p_{2}$ spans from 0 to $10^{-4}$. White contour lines indicate fidelity thresholds, illustrating the degradation of circuit fidelity as gate errors increase. These boundaries provide insight into the interplay between single- and two-qubit gate errors and their collective impact on algorithm performance. Coherence (i.e., $T_{1}$ and $T_{2}$) errors are excluded from this analysis, as the focus is solely on evaluating the impact of operational errors for design-oriented considerations.
4 Methods
4.1 Fidelity Estimation Algorithm
Building upon the theoretical findings about fidelity loss under depolarizing noise presented in Section 2.1, we present the fidelity estimation algorithm used in the previous results. This algorithm predicts the expected fidelity of a quantum circuit based on its structure and the calibration data of the quantum processor on which the circuit is being executed, under the assumption that each gate introduces depolarizing noise. The algorithm is summarized in Algorithm 1.
Algorithm 1 Fidelity Estimation
Quantum circuit circ, $p_{ent}\in[0,1]$
$F_{circ}$ Circuit Fidelity
[…]
$F_{q}\leftarrow 1\quad\forall q\in\mathcal{Q}$triangleright Initialize qubit-wise fidelity to 1
[…]
     $d\leftarrow 2^{dim(\texttt{op.qubits})}$
[…]
     if $d==2$ then $\triangleright$ Single-qubit gate
[…]
         $F_{q_{i}}\leftarrow(1-p)\cdot F_{q_{i}}+(1-p_{ent})\cdot\frac{p}{d}$
     else if $d==4$ then $\triangleright$ Two-qubit gate
[…]
         $\eta\leftarrow\frac{1}{2}\left(\sqrt{(1-p)\left(F_{q_{i}}+F_{q_{j}}\right)^{2}+p}-\sqrt{1-p}\left(F_{q_{i}}+F_{q_{j}}\right)\right)$
         $F_{q_{i}}\leftarrow\sqrt{1-p}\cdot F_{q_{i}}+(1-p_{ent})\cdot\eta$
         $F_{q_{j}}\leftarrow\sqrt{1-p}\cdot F_{q_{j}}+(1-p_{ent})\cdot\eta$
     end if
end for
return $\prod_{q\in\mathcal{Q}}F_{q}$
The calibration data provides gate fidelities for all single- and two-qubit gates ($d=2$ and $d=4$ respectively) of a given quantum processor. From this, we derive the depolarization parameter $p$ for each single- and two-qubit gate, ensuring that the gate fidelity $F_{g}$ from calibration aligns with the theoretical fidelity loss due to a depolarization factor $p$ (obtained in Corollary 1), which is given by:
$p=\frac{d(F_{g}-1)}{1-d}$ (39)
Leveraging the unitary invariance of both the depolarizing channel and quantum fidelity, all gates in the circuit are effectively replaced with fidelity-equivalent depolarizing channels. The proposed algorithm initializes each qubit with a fidelity of 1 and iteratively reduces the fidelity based on the depolarizing channels applied to each quantum state. The global fidelity is determined using the multiplicativity under tensor product property of fidelity. Specifically, the circuit’s overall fidelity ($F_{circ}$) is obtained as the product of the individual qubit-wise fidelities ($F_{q}$):
$F_{circ}=\prod_{q\in\mathcal{Q}}F_{q}$ (40)
As proved in Theorem 4, the fidelity loss induced by a depolarizing channel is influenced by the level of entanglement of the involved qubits. Accounting for this would require simulating all quantum states and tracking how each gate impacts the states throughout the circuit (i.e., to simulate the whole computation classically). This approach, however, is computationally prohibitive for systems with more than a few dozen qubits [27]. To address this, we introduce a hyperparameter $p_{ent}$, which defines the level of entanglement in the quantum states. No entanglement is assumed when $p_{ent}=0$, minimizing the fidelity loss and yielding an upper bound on the circuit’s fidelity. Conversely, setting $p_{ent}=1$ assumes maximum entanglement, resulting in a more significant fidelity loss for each gate and providing a lower bound on the circuit’s fidelity.
The results in Figure 3 reveal that the fidelity range determined by $p_{ent}=0$ and $p_{ent}=1$ is relatively narrow, in contrast to QVA, where the hyperparameter $w$ can cause fidelity estimates to vary by as much as 0.8. This demonstrates that an intermediate value, such as $p_{ent}=0.5$, can reasonably estimate the fidelity. Consequently, the inability to precisely determine the level of entanglement does not significantly impact the accuracy of the fidelity estimation.
To validate the approach presented in this work, it is essential to incorporate the effects of coherence errors [20, 3, 4] into the fidelity estimation model, as these errors are a significant source of decoherence in current quantum hardware. The exact decoherence induced by relaxation ($T_{1}$) and dephasing ($T_{2}$) depends on the particular quantum state (i.e., they are not unitary-invariant), which can only be fully assessed by simulating the quantum state. Instead, we assume a worst-case scenario where qubits undergo maximal decoherence, thus providing an upper bound on fidelity loss.
The resulting upper bound on fidelity loss is incorporated into Algorithm 1 by scaling the qubit-wise fidelity after the execution of each layer (i.e., a set of gates that can be applied simultaneously on independent qubits) of the circuit to account for coherence errors. Specifically, the relaxation time $T_{1}$ and dephasing time $T_{2}$ for each qubit are obtained from the device calibration data, and the execution time for each layer is estimated based on the duration of the gates that can be executed in parallel. After each layer, the fidelity of each qubit is updated as:
$\forall{q\in\mathcal{Q}}:F_{q}=F_{q}\cdot e^{-t_{\text{layer}}/T_{1}^{q}}\cdot\left(\frac{1}{2}e^{-t_{\text{layer}}/T_{2}^{q}}+0.5\right)$ (41)
This equation accounts for both relaxation and dephasing effects on fidelity, consistent with how IBMQ calibrates coherence times, modeling $T_{1}$ and $T_{2}$ using $A\cdot e^{-t/T}+B$. Resulting in $A\simeq 1$ and $B\simeq 0$ for $T_{1}$, and $A\simeq B\simeq 0.5$ for $T_{2}$ [47, 48].
4.2 Quantum Processor Execution
This section details the experimental setup and workflow for evaluating circuit success rates and comparing fidelity estimation methods. Figure 7 provides an overview of the entire workflow.
Figure 7: Workflow for obtaining fidelity predictions from the four estimation methods (Qiskit, ESP, QVA, proposed model, and proposed $+T_{1,2}$) and for determining the success rate from real-device executions. The process includes circuit selection, compilation to match device-specific constraints, fidelity estimation based on noise models and calibration data, execution on a real quantum processor, and measurement error mitigation to compute the success rate. This workflow ensures a consistent comparison of predictions with hardware performance.
We selected circuits ranging from 2 to 8 qubits ($q$) sourced from MQT Bench [49]. Each circuit $\mathcal{C}$ was concatenated with its inverse $\mathcal{C}^{-1}$, hence ensuring the final quantum state being equal to the initial state, the all-zero state $|0\rangle^{\otimes q}$ in our case, with a corresponding ideal measurement outcome of ’0’×$\times q$.
Before running the circuits on the actual quantum device, they went through the compilation process, which decomposed gates into native gates supported by the device [50, 51], mapped logical qubits to physical qubits [52], and added routing operations [53, 54] to meet hardware connectivity constraints [55]. We used Qiskit [36] for all stages of this compilation process, ensuring consistency across all the evaluated circuits. The result is a compiled circuit optimized for execution on the chosen quantum processor.
For each estimation method, fidelity predictions were made using the compiled circuit and calibration data from the quantum processor. For Qiskit simulations, noise models were directly obtained from the Qiskit library [36], tailored to the target device, and simulations were conducted using the density_matrix method to account for mixed-state evolution under noise.
Experiments were conducted on an IBM Eagle r3 processor [37] with 127 qubits. The compiler selected the most reliable physical qubits for each circuit. After execution, qubits were measured, and readout errors were mitigated using calibration data and assuming uncorrelated measurement errors between the qubits. The success rate is calculated as the ratio of correctly measured outcomes (’0’×$\times q$) over the total number of executions (2048 per circuit).
This setup establishes a robust framework for evaluating the performance of the real quantum device against fidelity predictions from various estimation methods, providing valuable insights into the accuracy and dependability of each model.
5 Data Availability
The datasets generated and/or analysed during the current study are available in the GitHub repository, https://github.com/escofetpau/Analytic-Model-of-Fidelity-under-Depolarizing-Noise.
6 Code Availability
The underlying code for this study is available in GitHub and can be accessed via this link https://github.com/escofetpau/Analytic-Model-of-Fidelity-under-Depolarizing-Noise.
7 Acknowledgements
This work was supported by the European Commission (QUADRATURE: 101099697, WINC: 101042080), MCIN and NextGenerationEU (QCOMM-CAT), Project PCI2022-133004 funded by MCIN/AEI/10.13039/501100011033, by the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda, and by the European Union NextGenerationEU/PRTR, and the ICREA Academia Award 2024. P.E. acknowledges support from an FPI-UPC grant funded by UPC and Banco Santander.
Appendix A Robustness of Fidelity Estimation Across Multiple Hardware Executions
Given the inherent variability in quantum hardware performance due to fluctuations in environmental conditions, calibration drift, and statistical noise in measurement processes, we conducted multiple independent experiments to validate the reliability of our fidelity estimation approach. Each experiment consisted of executing the same set of quantum circuits under comparable conditions, with device calibration data obtained prior to the execution to account for changes in hardware properties.
By repeating the experiments six times, we aimed to assess the consistency of our model’s predictions and its ability to generalize across different executions. The results confirm that our proposed approach systematically provides the most accurate fidelity estimates, demonstrating its robustness in real-world quantum computing scenarios. Figure 8 details the variations in fidelity predictions across experiments for all estimation methods, including coherence errors, and presents the corresponding $R^{2}$ scores to quantify the predictive accuracy in each case.
Figure 8: Comparison of the fidelity prediction accuracy for five different estimation models, each incorporating coherence errors ($T_{1}$ and $T_{2}$). The figure consists of five rows, one per model (including QVA with $w=0$), and six columns, each corresponding to an independent experiment conducted on different days using IBM Q Eagle processors (IBM Q Kyiv, IBM Q Sherbrooke, and IBM Q Brisbane). Each cell represents the fidelity error for a given model in a specific experiment. The $R^{2}$ metric is reported for each model, quantifying the correlation between predicted and measured fidelities. Across all experiments, the proposed model consistently achieves the highest $R^{2}$ values, ranging from $0.87$ to $0.94$, demonstrating superior predictive accuracy and robustness.